OpenStudy (anonymous):

Find all the zeroes of the equation. –3x4+ 27x2 + 1200 = 0

OpenStudy (anonymous):

-3(x^2+16)(x-5)(x+5) = 0 what are the zeros?

OpenStudy (anonymous):

Let x^2 = y nw -3x4+ 27x2 + 1200 = 0 -3y^2 + 27y + 1200 = 0 nw it is a quadratic equation use quad Formula to find zeros

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

-3(x^4 -9x^2-400)=0 or (x^4 - 25x^2 +16x^2 -400=0 or (x^2 -25 )(x^2+16)=0 hence (x+5)(x-5)(x+4i)(x-4i)=0 hence x=5,-5,4i,-4i (the required zeroes) here i=sqrt(-1)